OpenStudy (swagster23780):

Multiply. Express your answer in lowest terms. 3/8 x 2/7 = 7

Mehek (mehek14):

\(\dfrac{3*2}{8*7}=\)

OpenStudy (swagster23780):

Idk

Mehek (mehek14):

what is 3*2? 8*7?

OpenStudy (swagster23780):

3x2=6 and 8x7= 56

Mehek (mehek14):

\(\dfrac{6}{56}\) find the gcf of 6 and 56

OpenStudy (swagster23780):

3 and 2 for 6 and 8 and 7 for 56

OpenStudy (swagster23780):

@Mehek14

Mehek (mehek14):

no the ones that go in 6 and 56 like 2 \(\dfrac{6}{56}=\dfrac{3}{28}\)

OpenStudy (swagster23780):

So It Dividing

Mehek (mehek14):

yup divide the number that goes in both 6 and 56